# VIEWER: an extensible visual analytics framework for enhancing mental healthcare

### v1

@@ Results
VIEWER’s second use case was to facilitate medication review, a critical part of healthcare to ensure that patients receive the most appropriate and effective pharmacological treatments. Many individuals in secondary mental health services take psychiatric medications, often for long periods [40]. Regular reviews are essential for optimizing treatment outcomes, improving adherence to prescribed medications, and minimizing adverse effects [41]. This process requires a thorough assessment of all medication records within the entire EHR. Since most prescribing information and related contextual details (e.g. adverse reactions and non-adherence) are recorded in unstructured text such as clinical notes and correspondences between healthcare professionals (e.g. mental healthcare specialists, pharmacists, and GPs), each review typically takes hours to complete in the local Trust.
@@
By integrating data aggregation processes in data ingestion pipelines and filters within Kibana visualisations, VIEWER allows team members to access consistent information about their caseloads, providing a single source of truth. Also, interactive data visualisations in VIEWER enable users to complete these tasks with just a few clicks, significantly improving efficiency. For example, team leaders need to regularly review the number of patients under each care coordinator during supervision, a task that typically takes at least 20 minutes per review. In contrast, this can be accomplished with a single click of a filter for a team name or a care coordinator’s name in VIEWER. Additionally, VIEWER’s rapid development capability has enabled the Trust to meet CQC requirements and deliver results for an annual inspection within just 6 months.
@@
Usability and acceptability
Figure 5: The average response to each question in the System Usability Scale (SUS) survey, with a scale ranging from 1 (strongly disagree) to 5 (strongly agree).
To assess the usability of VIEWER, we used the standard System Usability Scale (SUS), which consists of a 10-item questionnaire with 5 response options ranging from “Strongly agree” to “Strongly disagree” [42] during its early development stage. 17 users completed the questionnaire, where 16 had used VIEWER, while 1 had not and gave mainly neutral responses (leading to their exclusion from the analysis). The overall score was 74/100, indicating “good” usability. Figure 5 presents the detailed scores for each SUS question from the 16 respondents. The positive scores are further supported by the respondents’ feedback in SI Table S2. Among the responses, 4 users scored the system over 80, indicating potential promoters of the system, while 2 users scored below 55, making them potential detractors. Both potential detractors provided free-text responses stating that they found the system useful and had used it to improve patient care but expressed a need for more training.
@@ Discussion
VIEWER also offers strong interoperability, allowing for directly deploying on top of existing EHR systems or incorporating into clinical research platforms to facilitate the translation of research into practice. In our implementation, we leveraged the clinical research infrastructure in our institution and use VIEWER to facilitate the transformation of research outcomes, particularly in data linkage and NLP pipelines within CRIS, into real-world clinical practices. As a result, the extensive set of clinical information extraction and NLP pipelines, developed and refined through multiple CRIS research projects over the last decade, has now been applied to inform clinical decisions directly at the point of care. This also enhances the accessibility of clinical research datasets and enables large-scale studies at the population level, e.g. examining outcome disparities across populations [43]. However, when a CRIS-like system is absent, CogStack, the backbone component used for data ingestion and visualisation in VIEWER, provides a set of open-source tools capable of parsing text content from binary documents (e.g. PDFs, scanned images, and Word files) and NLP packages for extracting medical information from clinical text [^5] https://github.com/CogStack/CogStack-Pipeline. Indeed, CogStack has been deployed at multiple NHS sites, including both mental and physical healthcare providers in the UK and beyond [30]. This offers a high degree of flexibility for easy customisation when integrating VIEWER with various EHR systems directly.
@@
Another notable strength of VIEWER is its remarkable extensibility and customizability. Although originally developed within mental healthcare settings, the VIEWER framework is vendor-agnostic and capable of importing data from various relational databases and external systems such as sensor devices and fitness monitors. While different settings have varying priorities and workflows, our data model design closely aligns with the Observational Medical Outcomes Partnership (OMOP) Common Data Model, an open community data standard for observational health data sciences and informatics [44], with modifications tailored to local settings and project-specific requirements. Also, our interface design adhere to well-established health informatics design methods [33, 44] and are general to cover key components in clinical practice. This versatility enables VIEWER to be effectively applied in other settings to enhance data usage and supporting informed clinical decision-making. Also, VIEWER supports extensive customisation options, including visualisation coding tools like Vega [45] and graphical user interfaces (GUIs), as well as integration with external libraries for predictive and statistical learning algorithms available in languages such as Python and R. This customizability empowers both technical and non-technical users to design and refine data visualisations and analytical dashboards. In fact, most visualisations in VIEWER were developed by clinicians rather than technicians, which has enhanced its clinical relevance and facilitated its rapid adoption in practice.
@@
    Interdisciplinary collaboration: As highlighted previous guidelines [46], the success of CDSS hinges on their ability to tackle real-world challenges faced by clinicians. To ensure this, it is essential to co-design a system with input from a broad range of stakeholders with interdisciplinary backgrounds, particularly clinicians and governance committees. From our experience in implementing VIEWER, this collaborative approach not only enhances the understanding of practical challenges and helps identify relevant solutions more effectively, but also fosters the involvement of key stakeholders as part of an iterative process to support the creation of practical models that are more likely to be adopted effectively.
@@
VIEWER and this study of its use also have limitations and requires further improvement. First, although a range of NLP applications have been used to extract useful information from clinical text, which contains the majority of patient information, this task remains challenging. The primary difficulties stem from the extensive use of medical terminology, grammatical inaccuracies, and the sensitivity of clinical data which results in limited availability of data for training large language models [52]. Our users have reported that temporality (distinguishing past, present, or future events), certainty (differentiating confirmed and hypothetical events), and negation are the most common sources of errors in NLP extraction results, highlighting areas that require further improvement [53]. As a contingency solution, VIEWER provides text snippets that offer complete context for each extracted event, allowing users to interpret and validate the NLP results effectively. Moreover, real-time extraction of information and aggregation of NLP-extracted information to construct care trajectories proves useful but also require further development. Second, beyond functionality, the sustainability of healthcare informatics solutions — encompassing aspects such as maintainability, scalability, and interoperability — is crucial for enhancing the quality and affordability of healthcare delivery. To fully understand the benefits and limitations of VIEWER, it is important to compare it with other open-source or commercial solutions that offer similar capabilities. A more extensive comparison, based on longer and broader usage, would provide valuable insights into whether the findings are widely applicable. Third, VIEWER has primarily been used by clinicians at SLaM. However, it would be beneficial to make VIEWER’s patient chart features available to individual patients, enabling them to monitor their own health metrics and input relevant health data. To explore this, we would need authorisation to grant patient access to the system, as well as technical support and data integration services to connect with other patient portals. Finally, VIEWER has only been deployed at SLaM so far and thus our evaluation focused on findings in a single institution. However, as VIEWER was designed to be EHR-agnostic, porting it to another EHR system is feasible. We are eager to deploy VIEWER in other institutions and assess the generalizability of our findings. Moreover, our current evaluation relies on observational research methods. To formally assess the effectiveness of our informatics intervention and determine the factors contributing to its success across various teams and settings, randomised controlled trials are needed.
Conclusion
@@
The authors also would like to express their gratitude to Damian Larkin, Ian Cutting-Jones, Omar Rayner-Andrews, Percy Chinouya, Juliet Hurn, Abigail Bennett, Justine Chow, Man Yuen, Shubhra Mace, Sandra Batchelor, Isaac Enakimio, and many others for their valuable advice and contributions to the project.

### v2

@@ 3.4.2 Medication review
VIEWER’s second use case was to facilitate medication review for 164,819 patients with documented prescriptions for psychiatric medications, including antipsychotics, antidepressants, and mood stabilisers. Many individuals in secondary mental health services take psychiatric medications, often for long periods [46]. Regular reviews are essential for optimizing treatment outcomes, improving adherence to prescribed medications, and minimizing adverse effects [47]. This process requires a thorough assessment of all medication records within the entire EHR. Since most prescribing information and related contextual details (e.g. adverse reactions and non-adherence) are recorded in unstructured text such as clinical notes and correspondences between healthcare professionals (e.g. mental healthcare specialists, pharmacists, and GPs), each review typically takes hours to complete in the local Trust.
@@ 3.4.3 Caseload management
By integrating data aggregation processes in data ingestion pipelines and filters within Kibana visualisations, VIEWER allows team members to access consistent and unified information about their caseloads. Also, interactive data visualisations in VIEWER enable users to complete these tasks with just a few clicks, significantly improving efficiency. For example, team leaders need to regularly review the number of patients under each care coordinator during supervision, a task that typically takes at least 20 minutes per review. In contrast, this can be accomplished with a single click of a filter for a team name or a care coordinator’s name in VIEWER. Additionally, VIEWER’s rapid development capability has enabled the Trust to meet CQC requirements and deliver results for an annual inspection within just 6 months.
@@
3.5 Usability and acceptability
Figure 5: Proportions of various responses to each question in the SUS survey.
To assess the usability of VIEWER, we used the standard System Usability Scale (SUS), which consists of a 10-item questionnaire with 5 response options ranging from “Strongly agree” to “Strongly disagree” [48] during its early development stage. 16 users completed the questionnaire, and the overall score was 74/100, indicating “good” usability. Figure 5 presents the detailed responses for each SUS question from the 16 respondents. The positive scores are further supported by the respondents’ feedback in SI Table S2. Among the responses, 4 users scored the system over 80, indicating potential promoters of the system, while 2 users scored below 55, making them potential detractors. Both potential detractors provided free-text responses stating that they found the system useful and had used it to improve patient care but expressed a need for more training.
@@ 4 Discussion
VIEWER also offers high potential for interoperability, allowing for directly deploying on top of existing EHR systems or incorporating into clinical research platforms to facilitate the translation of research into practice. In our implementation, we leveraged the clinical research infrastructure in our institution and use VIEWER to facilitate the transformation of research outcomes, particularly in data linkage and NLP pipelines within CRIS, into real-world clinical practices. As a result, the extensive set of clinical information extraction and NLP pipelines, developed and refined through multiple CRIS research projects over the last decade, has now been applied to inform clinical decisions directly at the point of care. This also enhances the accessibility of clinical research datasets and enables large-scale studies at the population level, e.g. examining outcome disparities across populations [35]. However, when a CRIS-like system is absent, CogStack, the backbone component used for data ingestion and visualisation in VIEWER, provides a set of open-source tools capable of parsing text content from binary documents (e.g. PDFs, scanned images, and Word files) and NLP packages for extracting medical information from clinical text [^6] https://github.com/CogStack/CogStack-Pipeline. Indeed, CogStack has been deployed at multiple NHS sites, including both mental and physical healthcare providers in the UK and beyond [31]. This offers a high degree of flexibility for easy customisation when integrating VIEWER with various EHR systems directly.
@@
Another notable strength of VIEWER is its extensibility and customizability. Although originally developed within mental healthcare settings, the VIEWER framework is vendor-agnostic and capable of importing data from various relational databases and external systems such as sensor devices and fitness monitors. While different settings have varying priorities and workflows, our data model design closely aligns with the Observational Medical Outcomes Partnership (OMOP) Common Data Model, an open community data standard for observational health data sciences and informatics [49], with modifications tailored to local settings and project-specific requirements. Also, our interface design adhere to well-established health informatics design methods [34, 49] and are general to cover key components in clinical practice. This versatility enables VIEWER to be effectively applied in other settings to enhance data usage and supporting informed clinical decision-making. Also, VIEWER supports extensive customisation options, including visualisation coding tools like Vega [50] and graphical user interfaces (GUIs), as well as integration with external libraries for predictive and statistical learning algorithms available in languages such as Python and R. This customizability empowers both technical and non-technical users to design and refine data visualisations and analytical dashboards. In fact, most visualisations in VIEWER were developed by clinicians rather than technicians, which has enhanced its clinical relevance and facilitated its rapid adoption in practice.
@@
    Interdisciplinary collaboration: As highlighted previous guidelines [51], the success of CDSS hinges on their ability to tackle real-world challenges faced by clinicians. To ensure this, it is essential to co-design a system with input from a broad range of partners with interdisciplinary backgrounds, particularly clinicians and governance committees. From our experience in implementing VIEWER, this collaborative approach not only enhances the understanding of practical challenges and helps identify relevant solutions more effectively, but also fosters the involvement of key partners as part of an iterative process to support the creation of practical models that are more likely to be adopted effectively.
@@
This includes incorporating robust privacy-preserving techniques, such as differential privacy, and ensuring that all stakeholders (clinicians, patients, and developers) are properly informed about how their data is being used and processed. Ethical considerations, especially regarding mental health data, will be crucial as we continue to develop and scale such technologies.
@@
VIEWER and this study of its use also have limitations and requires further improvement. First, although a range of NLP applications have been used to extract useful information from clinical text, which contains the majority of patient information, this task remains challenging. The primary difficulties stem from the extensive use of medical terminology, grammatical inaccuracies, and the sensitivity of clinical data which results in limited availability of data for training large language models [57]. Our users have reported that temporality (distinguishing past, present, or future events), certainty (differentiating confirmed and hypothetical events), and negation are the most common sources of errors in NLP extraction results, highlighting areas that require further improvement [58]. As a contingency solution, VIEWER provides text snippets that offer complete context for each extracted event, allowing users to interpret and validate the NLP results effectively. Moreover, real-time extraction of information and aggregation of NLP-extracted information to construct care trajectories proves useful but also require further development. Given the sensitive nature of mental healthcare data, it is essential to prioritize patient confidentiality and privacy in the development of these technologies. This requires strict adherence to ethical guidelines, such as data anonymization, and compliance with regulations like the GDPR and the NHS Data Security and Protection Toolkit.
Second, beyond functionality, the sustainability of healthcare informatics solutions—encompassing aspects such as maintainability, scalability, and interoperability—is crucial for enhancing the quality and affordability of healthcare delivery. To fully understand the benefits and limitations of VIEWER, it is important to compare it with other open-source or commercial solutions that offer similar capabilities. A more extensive comparison, based on longer and broader usage, would provide valuable insights into whether the findings are widely applicable. Third, VIEWER has primarily been used by clinicians at SLaM. However, it would be beneficial to make VIEWER’s patient chart features available to individual patients, enabling them to monitor their own health metrics and input relevant health data. To explore this, we would need authorisation to grant patient access to the system, as well as technical support and data integration services to connect with other patient portals.
Finally, VIEWER has only been deployed at SLaM so far and thus our evaluation focused on findings in a single institution. However, as VIEWER was designed to be EHR-agnostic, porting it to another EHR system is feasible. We are eager to deploy VIEWER in other institutions and assess the generalizability of our findings. Moreover, this paper has primarily focused on outlining the technical components and development process of VIEWER, and our current evaluation relies on observational research methods. To formally assess the effectiveness of our informatics intervention and determine the factors contributing to its success, such as the types of use cases and participant characteristics, including backgrounds, demographics, and areas of expertise, across various teams and settings, randomised controlled trials are needed.
5 Conclusion
@@
Competing Interests Statement
RS declares research support received in the last 36 months from Janssen, GSK and Takeda, and royalties from Oxford University Press.
Contributorship Statement
TW, DC, YJM, MB, and RH contributed to the conceptualization, methodology design, data acquisition, analysis, interpretation, and software development. Data extraction and preprocessing were carried out by MB, TW, YJM, and DK, with guidance from DC and RH. Method development was led by TW, DC, and YJM under the guidance of CP, CDP, TS, BA, SM, ZK, MD, AR, RS, PM, RD, and RH. Funding acquisition and supervision were supported by AR, RS, PM, and RD. TW drafted the initial manuscript and implemented revisions, while all authors participated in reviewing and editing the final version.
Acknowledgements
The authors would like to express their gratitude to Damian Larkin, Ian Cutting-Jones, Omar Rayner-Andrews, Percy Chinouya, Juliet Hurn, Abigail Bennett, Justine Chow, Man Yuen, Shubhra Mace, Sandra Batchelor, Isaac Enakimio, and many others for their valuable advice and contributions to the project.
Data Availability
All data utilized in this study are derived from patient records within SLaM and are not publicly accessible. Details about the software and tools employed in the research are provided within the article, including access links where available or upon request.